A solution is made by mixing 41.0 mL of ethanol, C2H6O, and 59.0 mL of water. Assuming ideal behavior, what is the vapor pressure of the solution at 20 °C? Values at 20 C: ethanol, density = 0.789 g/mL, VP = 43.9 torr water, density = 0.998 g/ml, VP = 17.5 torr
Answer #1

moles of ethanol: 41.0 mL x 0.789 g/mL x 1/(46.068 g/mol) = 0.7022 mol

mole of water: 59.0 mL x 1.0 g/mL x 1/(18.02 g/mol) = 3.274 mol

total number of moles: 3.9762

mole fraction of ethanol: 0.7022/3.9762 = 0.1766

mole fraction of water: 0.8234

VPtot = mole fraction ethanol x VPethanol + mole fraction water x VPwater

VPtot = (0.1766)x(43.9 torr) + (0.8234)x(17.5 torr) = 22.1622 torr

hope this helped!